OpenStudy (anonymous):

If y is directly proportional to the square root of x, and y = 25 when x = 3, find x when y = 100.

OpenStudy (anonymous):

√x/100 = √3/25 √x = (√3/25)(100) x = [(√3/25)(100)]^2 x = 48 Hence, x = 48 when y = 100 does that help?

OpenStudy (anonymous):

yes it does thnks :)

OpenStudy (anonymous):

welcome:)

OpenStudy (anonymous):

can u help me w another one?

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

The volume of a cone varies jointly as the height and the square of the radius of the base. A cone height 8 cm and a base of diameter 9 cm has volume 54π. Find the constant of variation and use it to develop the general formula for the volume of a cone.

OpenStudy (anonymous):

is it multiple chose

OpenStudy (anonymous):

yes V = 0.33π , V = 54πr2h V = 0.33πr2h, V = 3πr

OpenStudy (anonymous):

V = (1/3)h*r^2 this is what i came up with

OpenStudy (anonymous):

im not sure how to turn in 1/3 into a decimal

OpenStudy (anonymous):

v= 0.33 pi r2h?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

aa okaii :)